# On the capability of high redshift kSZ measurement with galaxy surveys

### v1

@@ Appendix C The equivalence between kSZ measurement methods
Figure 6: The correlation coefficient of different cluster mass bins (left) and the correlation function (right) of the true and reconstructed velocity fields in three redshift bins $1.00.6<z<$ 1.0, $1.61.0<z<$ 1.6 and $2.41.6<z<$ 2.4 (from top to bottom). The black lines show the case that reconstructing velocity field directly from $\delta_{g}$ without applying any filters. The red lines show the case of applying the Wiener filter using the main text. The dashes lines show the case of applying the Gaussian filter with different radius $R_{g}$.
@@ Appendix D Test the velocity reconstruction method
In this section, our research delves into the impact of applying different filters to the density field on velocity reconstruction performance. In the previous literatures, Gaussian filters in Fourier space, represented as $W_{g}(k)=\exp(-k^{2}R_{s}^{2})$, have been utilized to mitigate shot noise in galaxy density fields, with $R_{s}$ being a pivotal, adjustable parameter for optimizing velocity reconstruction. In figure 6, we present a comparative analysis of the correlation coefficients across various mass bins of clusters and the correlation function between true and reconstructed velocities. Our findings suggest that, in the stacking method, the performance of the optimal Gaussian filter is comparable to the Wiener filter discussed in the main text. However, the Wiener filter demonstrates a noticeable advantage in the tomography method at smaller scales $k>0.1[h/{\rm Mpc}]$, highlighting the Wiener filter’s efficacy in enhancing velocity reconstruction at these scales.

### v2

@@ Appendix C The equivalence between kSZ measurement methods
Figure 7: The correlation coefficient of different galaxy group mass bins (left) and the correlation function (right) of the true and reconstructed velocity fields in three redshift bins $1.00.6<z<$ 1.0, $1.61.0<z<$ 1.6 and $2.41.6<z<$ 2.4 (from top to bottom). The black lines show the case that reconstructing velocity field directly from $\delta_{g}$ without applying any filters. The red lines show the case of applying the Wiener filter using the main text. The dashes lines show the case of applying the Gaussian filter with different radius $R_{g}$.
@@ Appendix D Test the velocity reconstruction method
In this section, our research delves into the impact of applying different filters to the density field on velocity reconstruction performance. In the previous literatures, Gaussian filters in Fourier space, represented as $W_{g}(k)=\exp(-k^{2}R_{s}^{2})$, have been utilized to mitigate shot noise in galaxy density fields, with $R_{s}$ being a pivotal, adjustable parameter for optimizing velocity reconstruction. In figure 7, we present a comparative analysis of the correlation coefficients across various mass bins of galaxy groups and the correlation function between true and reconstructed velocities. Our findings suggest that, in the stacking method, the performance of the optimal Gaussian filter is comparable to the Wiener filter discussed in the main text. However, the Wiener filter demonstrates a noticeable advantage in the tomography method at smaller scales $k>0.1[h/{\rm Mpc}]$, highlighting the Wiener filter’s efficacy in enhancing velocity reconstruction at these scales.